OpenStudy (anonymous):

A helium balloon contains 0.48g of helium. Given the molecular weight of helium is 4.0 g/mol, how many moles of helium does it contain?

OpenStudy (accessdenied):

The unit analysis works as such: \( \rm grams \times \dfrac{moles}{gram} = grams / \left( \dfrac{grams}{mole} \right) = moles \) So we want to take our mass in grams and divide by the molecular weight grams/mole. 0.48 g / 4.0 g/mol = ? mol